# Building Efficient Universal Classifiers with Natural Language Inference

### 29 Dec 2023 v1

@@ 2 NLI as a universal task
The Natural Language Inference (NLI) task [^2] An older but more expressive name for the task is RTE, Recognising Textual Entailment (Dagan et al., 2006) is defined as recognising if the meaning of one text (the hypothesis) is entailed in another text (the premise). For example, the hypothesis “The EU is not trustworthy" is entailed in the premise “The EU has betrayed its partners during the negotiations on Sunday". To create NLI datasets, workers are presented with a text (the premise) and are tasked with writing a hypothesis that is either clearly true given the premise (entailment), clearly false given the premise (contradiction), or that might be true or false but is not clearly entailed or a contradiction (neutral). Several large scale NLI datasets with hundreds of thousands of unique hypothesis-premise pairs for these three classes have been created by crowd workers or language models (Bowman et al., 2015; Williams et al., 2018; Conneau et al., 2018; Nie et al., 2020; Parrish et al., 2021; Liu et al., 2022). For simplicity and to increase universality, the task can be simplified into a binary entailment vs. not-entailment task by merging the ’neutral’ and ’contradiction’ labels (Yin et al., 2021).
@@
This binary NLI task is universal, because any text classification task can be reformulated into this entailment vs. not-entailment decision through label verbalisation (see figure 1). Take topic classification as an example. The task could be to determine if the text “We need to raise tariffs" belongs to the topic “economy" or “welfare". From an NLI perspective, we can interpret the text “We need to raise tariffs" as the premise and verbalise the topic labels in two topic hypotheses: “This text is about economy" and “This text is about welfare". The classification task reformulated as an NLI task then consists in determining which of the two topic hypotheses is more entailed in the text of interest (premise). In different words: Which hypothesis is more consistent with the text of interest?
A model fine-tuned on NLI data (e.g. “BERT-NLI") can then be used to test any hypotheses formulated by a human against any text of interest (premise). For each individual hypothesis-premise pair, an NLI models will output a probability for entailment and not-entailment. To choose the most probable topic, we can select the hypothesis with the highest entailment score. Following the same procedure, any other text classification task can be reformulated as an NLI task, from stance detection, sentiment classification to factuality classification (see figure 1). Any class can be verbalised as a hypothesis (similar to the prompt of a generative LLM) and can then be tested against any text of interest. [^3] Note that an NLI model will always only do one task (NLI) just like a GPT model can only predict the next token. These tasks are universal because any other specific task can be reformatted into these more general tasks.
The main disadvantage of NLI for universal classification is that it requires a separate prediction for each of N class hypotheses, creating computational overhead for tasks with many classes. The main advantage is that identifying a new class only requires verbalising it as a hypotheses and passing it to an NLI model without the need of fine-tuning a new task-specific model from scratch (zeroshot classification). The most prominent implementation of this approach is probably the Hugging Face ZeroShotClassificationPipeline (see figure 2) which uses this NLI-based approach under the hood (Wolf et al., 2020). [^4] https://huggingface.co/docs/transformers/v4.21.2/en/main_classes/pipelines##transformers.ZeroShotClassificationPipeline The models created in the paper are designed to be directly compatible with this pipeline.
@@ 3 A guide to building a universal classifier
In this guide we explain how this type of universal classifier is built. Each step is accompanied by a Jupyter notebook that implements each step end-to-end. [^5] https://github.com/MoritzLaurer/zeroshot-classifier The main steps are:
@@ 3.1 Data selection, preprocessing and harmonization
The resulting merged ~885000 hypothesis-premise pairs would be enough to train a decent NLI model capable of zeroshot classification. The NLI datasets were, however, not created with zeroshot classification in mind. Crowd workers were instructed to write hypotheses the are entailed, contradictory or neutral towards a text, which led to a wide range of hypothesis-premise pairs. They were not specifically instructed to create data for typical classification tasks such as identifying topics, sentiment, stances, emotions, toxicity, factuality etc. which users might be interested in in practice (e.g. “This text is about topic X"). To improve performance on these types of tasks, we therefore add a second collection of standard non-NLI classification datasets reformatted into the NLI format.
@@
We choose 28 popular non-NLI datasets with diverse classification tasks linked to sentiment, emotions, intent, toxicity, bias, topics, factuality, spam etc. with 387 classes in total. We selected most datasets based on their popularity (downloads) on the Hugging Face Hub. We also add some non-NLI datasets that are not available on the Hugging Face hub and create separate preprocessing notebooks for each of them (e.g. 1-data-harmonization-manifesto.ipynb). The full list of datasets with information on tasks, licenses and data quality is available in our dataset overview file. [^7] https://github.com/MoritzLaurer/zeroshot-classifier/blob/main/datasets_overview.csv
@@ 3.3 Hypothesis formulation and NLI formatting
For each row in each non-NLI training dataset we now add a new “hypothesis" column with the correct class hypotheses corresponding to the respective text. Moreover, in a new “label" column, these text-hypothesis pairs receive the label “0" for “entailment". We then multiply each text by two and pair the copied text with a random incorrect class hypotheses and the label “1" for “not-entailment". This multiplication ensures that the model does not only learn that class hypotheses are always true and it functions as a form of data augmentation. When we rename the “text" column to “premise", this dataset now has exactly the same format as the NLI dataset with the columns “premise", “hypothesis", “label" for binary entailment vs. not-entailment classification. This conversion is implemented in the function format_nli_trainset. We can now simply concatenate the non-NLI and the NLI training data.
@@ 5 Limitations
We outline several limitations of this paper and invite readers to improve on our implementation. First, while we have included 28 non-NLI datasets, the diversity of these academic datasets is limited and they do not cover the full diversity of classification use-cases users will need in practice. The instruction fine-tuning literature for generative LLMs has shown the potential of using SotA models like GPT-4 to generate diverse training data and distilling their capabilities into much smaller models (Taori et al., 2023; Tunstall et al., 2023). While many such datasets exist for generative tasks, hardly any are available for encoder-only classifiers like BERT (Sileo, 2023; Longpre et al., 2023a, b). We assume that smart LLM prompting could result in a more diverse dataset than our collection and could further improve generalisation.
Second, we assume that our data still contains a certain degree of noise. Additional data cleaning techniques could be used, for example discarding training data where the DeBERTa-v3 model still disagrees with the label after fine-tuning or targeted manual inspection enabled by active learning.
Third, an inherent limitation of NLI for zeroshot classification is that each additional class requires an additional forward pass (prediction) through the model. This makes the approach less suitable for tasks with a high amount of classes. At the same time, even if multiple forward passes are required, encoder-only models with only around a hundred million parameters are still more efficient than decoder models with multiple billion parameters while possibly being more accurate Xu et al. (2023); Schick and Schütze (2021b).
Fourth, we use the relatively old DeBERTa-v3 from November 2021 (He et al., 2021), which misses relevant recent innovations like longer context windows or flash attention (Dao et al., 2022). Unfortunately we are not aware of a better encoder-only model and releases have recently been dominated by larger generative decoder models.
Fifth, several other universal classification approaches exist that were beyond the scope of this paper: PET, which combines masked-language-modeling and label verbalisation (Schick and Schütze, 2021a), replaced-token-detection combined with prompts (Xia et al., 2022; Yao et al., 2022; Xu et al., 2023), question-answering (Bragg et al., 2021), or next-sentence-prediction as an interesting self-supervised alternative to NLI (Ma et al., 2021; Sun et al., 2022).

### 22 Mar 2024 v2

@@ 2 NLI as a universal task
The Natural Language Inference (NLI) task [^2] An older but more expressive name for the task is RTE, Recognising Textual Entailment (Dagan et al., 2006) is defined as recognising if the meaning of one text (the hypothesis) is entailed in another text (the premise). For example, the hypothesis “The EU is not trustworthy" is entailed in the premise “The EU has betrayed its partners during the negotiations on Sunday". To create NLI datasets, workers are presented with a text (the premise) and are tasked with writing a hypothesis that is either clearly true given the premise (entailment), clearly false given the premise (contradiction), or that might be true or false but is not clearly entailed or a contradiction (neutral). Several large scale NLI datasets with hundreds of thousands of unique hypothesis-premise pairs for these three classes have been created by crowd workers or language models (Bowman et al., 2015; Williams et al., 2018; Conneau et al., 2018; Nie et al., 2020; Parrish et al., 2021; Liu et al., 2022). For simplicity and to increase universality, the task can be simplified into a binary entailment vs. not-entailment task by merging the ‘neutral’ and ‘contradiction’ labels (Yin et al., 2021).
@@
This binary NLI task is universal, because any text classification task can be reformulated into this entailment vs. not-entailment decision through label verbalisation (see figure 1). Take topic classification as an example. The task could be to determine if the text “We need to raise tariffs" belongs to the topic “economy" or “welfare". From an NLI perspective, we can interpret the text “We need to raise tariffs" as the premise and verbalise the topic labels in two topic hypotheses: “This text is about economy" and “This text is about welfare". The classification task reformulated as an NLI task then consists of determining which of the two topic hypotheses is more entailed in the text of interest (premise). In different words: Which hypothesis is more consistent with the text of interest?
A model fine-tuned on NLI data (e.g. “BERT-NLI") can then be used to test any hypothesis formulated by a human against any text of interest (premise). For each individual hypothesis-premise pair, an NLI models will output a probability for entailment and not-entailment. To choose the most probable topic, we can select the hypothesis with the highest entailment score. Following the same procedure, any other text classification task can be reformulated as an NLI task, from stance detection, sentiment classification to factuality classification (see figure 1). Any class can be verbalised as a hypothesis (similar to the prompt of a generative LLM) and can then be tested against any text of interest. [^3] Note that an NLI model will always only do one task (NLI) just like a GPT model can only predict the next token. These tasks are universal because any other specific task can be reformatted into these more general tasks.
The main disadvantage of NLI for universal classification is that it requires a separate prediction for each of N class hypotheses, creating computational overhead for tasks with many classes. The main advantage is that identifying a new class only requires verbalising it as a hypothesis and passing it to an NLI model without the need of fine-tuning a new task-specific model from scratch (zeroshot classification). The most prominent implementation of this approach is probably the Hugging Face ZeroShotClassificationPipeline (see figure 2) which uses this NLI-based approach under the hood (Wolf et al., 2020). [^4] https://huggingface.co/docs/transformers/v4.21.2/en/main_classes/pipelines##transformers.ZeroShotClassificationPipeline The models created in the paper are designed to be directly compatible with this pipeline.
@@ 3 A guide to building a universal classifier
In this guide we explain how this type of universal classifier is built. Each step is accompanied by a Jupyter notebook available on GitHub that implements each step end-to-end. [^5] https://github.com/MoritzLaurer/zeroshot-classifier The main steps are:
@@ 3.1 Data selection, preprocessing and harmonization
The resulting merged ~885000 hypothesis-premise pairs would be enough to train a decent NLI model capable of zeroshot classification. The NLI datasets were, however, not created with zeroshot classification in mind. Crowd workers were instructed to write hypotheses that are entailed, contradictory or neutral towards a text, which led to a wide range of hypothesis-premise pairs. They were not specifically instructed to create data for typical classification tasks such as identifying topics, sentiment, stances, emotions, toxicity, factuality etc. which users might be interested in in practice (e.g. “This text is about topic X"). To improve performance on these types of tasks, we therefore add a second collection of standard non-NLI classification datasets reformatted into the NLI format.
@@
We choose 28 popular non-NLI datasets with diverse classification tasks linked to sentiment, emotions, intent, toxicity, bias, topics, factuality, spam etc. with 387 classes in total. We selected most datasets based on their popularity (downloads) on the Hugging Face Hub. We also add some non-NLI datasets that are not available on the Hugging Face hub and create separate preprocessing notebooks for each of them (e.g. 1-data-harmonization-manifesto.ipynb). The full list of datasets with information on tasks, licenses and data quality is available in our dataset overview file. [^7] https://github.com/MoritzLaurer/zeroshot-classifier/blob/main/v1_human_data/datasets_overview.csv
@@ 3.3 Hypothesis formulation and NLI formatting
For each row in each non-NLI training dataset we now add a new “hypothesis" column with the correct class hypothesis corresponding to the respective text. Moreover, in a new “label" column, these text-hypothesis pairs receive the label “0" for “entailment". We then multiply each text by two and pair the copied text with a random incorrect class hypothesis and the label “1" for “not-entailment". This multiplication ensures that the model does not only learn that class hypotheses are always true and it functions as a form of data augmentation. When we rename the “text" column to “premise", this dataset now has exactly the same format as the NLI dataset with the columns “premise", “hypothesis", “label" for binary entailment vs. not-entailment classification. This conversion is implemented in the function format_nli_trainset. We can now simply concatenate the non-NLI and the NLI training data.
@@ 5 Limitations
We outline several limitations of this paper and invite readers to improve on our implementation. First, while we have included 28 non-NLI datasets, the diversity of these academic datasets is limited and they do not cover the full diversity of classification use-cases users will need in practice. All datasets are only in English. The instruction fine-tuning literature for generative LLMs has shown the potential of using SotA models like GPT-4 to generate diverse training data and distilling their capabilities into much smaller models (Taori et al., 2023; Tunstall et al., 2023). While many such datasets exist for generative tasks, hardly any are available for encoder-only classifiers like BERT (Sileo, 2023; Longpre et al., 2023a, b). We assume that smart LLM prompting could result in a more diverse dataset than our collection and could further improve generalisation.
Second, the model comparisons are limited as we only compare BERT-NLI models among each other. We do not compare classification performance, inference speed, memory requirements, and costs to larger generative LLMs or APIs.
Third, we assume that our data still contains a certain degree of noise. Additional data cleaning techniques could be used, for example discarding training data where the DeBERTa-v3 model still disagrees with the label after fine-tuning or targeted manual inspection enabled by active learning.
Fourth, an inherent limitation of NLI for zeroshot classification is that each additional class requires an additional forward pass (prediction) through the model. This makes the approach less suitable for tasks with a high amount of classes. At the same time, even if multiple forward passes are required, encoder-only models with only around a hundred million parameters are still more efficient than decoder models with multiple billion parameters while possibly being more accurate (Xu et al., 2023; Schick and Schütze, 2021b).
Fifth, we use the relatively old DeBERTa-v3 from November 2021 (He et al., 2021), which misses relevant recent innovations like longer context windows or flash attention (Dao et al., 2022). Unfortunately we are not aware of a better encoder-only model and releases have recently been dominated by larger generative decoder models.
Sixth, several other universal classification approaches exist that were beyond the scope of this paper: PET, which combines masked-language-modeling and label verbalisation (Schick and Schütze, 2021a), replaced-token-detection combined with prompts (Xia et al., 2022; Yao et al., 2022; Xu et al., 2023), question-answering (Bragg et al., 2021), or next-sentence-prediction as an interesting self-supervised alternative to NLI (Ma et al., 2021; Sun et al., 2022).
@@ Appendix B Hypotheses per task
Appendix C Datasets
For details on all datasets used, see the overview table. [^13] https://github.com/MoritzLaurer/zeroshot-classifier/blob/main/v1_human_data/datasets_overview.csv To give citation credit to the authors of all datasets, here is the full list of dataset sources: Grano et al. (2017); Davidson et al. (2017); Saravia et al. (2018); Zhang et al. (2015); Almeida et al. (2011); Casanueva et al. (2020); Malo et al. (2014); Mathew et al. (2021); McAuley and Leskovec (2013); Soups (2015); Faruqui and Das (2018); Maas et al. (2011); FitzGerald et al. (2023); Pang and Lee (2005); Chatterjee et al. (2019); Sap et al. (2020); Rashkin et al. (2019); Adams et al. (2017); Gekhman et al. (2023); Unknown (2024); Parrish et al. (2021); Nie et al. (2020); Williams et al. (2018); Liu et al. (2022); Burst et al. (2020); Project (2015); Thorne et al. (2018)